OpenStudy (anonymous):

erica bought 3 1/2 yards of fabric, if she used 2 3rd of fabric. how much she has left?

OpenStudy (ivettef365):

you need to subtract 3 1/2 - 2/3

OpenStudy (ivettef365):

to subtract convert mixed number into fraction so you hvae 3 1/2 = 7/2 7 2 --- - --- 2 3

OpenStudy (anonymous):

\[3 \frac{1 }{ 2 }-\frac{ 2 }{ 3 }\]\[\frac{ 1 }{2 } \to \frac{ 3 }{ 6 } and \frac{ 2 }{ 3 } \to \frac{ 4 }{ 6 }\] Change

OpenStudy (anonymous):

\[2 \frac{ 3 }{ 6 }-\frac{ 4 }{ 6 }\]

OpenStudy (anonymous):

Can you figue it out from there?